OpenStudy (hdrager):

A hypothetical square grows at a rate of 25 m^2/min. How fast are the sides of the square increasing when the sides are 9 m. each?

OpenStudy (hdrager):

It's a related rates problem

OpenStudy (danjs):

Start with listing what you given...

OpenStudy (hdrager):

k so dA/dt= 25

OpenStudy (hdrager):

A=s^2

OpenStudy (danjs):

right, square with sides 's', now take the derivative of the area equation with respect to time t

OpenStudy (hdrager):

A=2s

OpenStudy (danjs):

remember the chain rule, you are differentiating both sides with respect to time t. \[\large A=s^2\]

OpenStudy (danjs):

\[\large \frac{ dA }{ dt }=\frac{ d }{ ds }[s^2]*\frac{ ds }{ dt }\]

OpenStudy (danjs):

\[\large \frac{ dA }{ dt }=2s*\frac{ ds }{ dt }\]

OpenStudy (danjs):

good with that?

OpenStudy (hdrager):

would it be 7?

OpenStudy (hdrager):

oops nvm

OpenStudy (hdrager):

\[\frac{ ds }{ dt }=\frac{ 25 }{ 18 }\]

OpenStudy (danjs):

yeah, you can use the differential equation to solve like that, they say the area stays at a rate of 25 m^2/min \[\large 25=2s* \frac{ ds }{ dt }\] at any time you have side length s, and a side rate of change of ds/dt

OpenStudy (danjs):

any other probs?

OpenStudy (hdrager):

not with this one. Thanks for walking me through it!

OpenStudy (danjs):

welcome